# On the Smallest Size of Internal Collage Systems

### v1

@@ 4 MAX-SAT formulation to compute $\hat{c}(T)$
  • $\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}$ for $(i^{\prime},i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow\circ,\circ}\rrbracket$ with $\ell^{\prime}=\ell$, $(i^{\prime},i)\in\llbracket\mathit{ref}^{\mathrm{B}}_{\circ\leftarrow\circ,\cdot}\rrbracket$ with $\ell^{\prime}=i-i^{\prime}$, or $(i^{\prime},\ell^{\prime},i)\in\llbracket\mathit{ref}^{\mathrm{C}}_{\circ,\circ\leftarrow\circ,\cdot}\rrbracket$: $\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}=1$ iff there is an ICS-factor that starts at $i$ and refers to $T[i^{\prime}..i^{\prime}+\ell^{\prime})$.
  • $\mathit{q}_{i^{\prime},\ell^{\prime}}$ for $(i^{\prime},\ell^{\prime})\in\llbracket\mathit{dref}_{\circ,\circ\leftarrow\cdot}\rrbracket$ or $(i^{\prime},i,\ell)\in\llbracket\mathit{ref}^{\mathrm{B}}_{\circ\leftarrow\circ,\cdot}\rrbracket$ with $\ell^{\prime}=\ell+i-i^{\prime}$: $\mathit{q}_{i^{\prime},\ell^{\prime}}=1$ iff $i^{\prime}..i^{\prime}+\ell^{\prime})$ is in the set $I$ defined in Lemma˜5.
@@
$\forall(i,\ell)\in\llbracket\mathit{f}_{\circ,\circ}\rrbracket\setminus(\llbracket\mathit{ref}^{\mathrm{A}}_{\cdot\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{B}}_{\cdot\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{C}}_{\cdot,\cdot\leftarrow\circ,\circ}\rrbracket)\mbox{ with }l>1:\lnot\mathit{f}_{i,l}$ (2)
@@
$\forall(i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\cdot\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{B}}_{\cdot\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{C}}_{\cdot,\cdot\leftarrow\circ,\circ}\rrbracket:\mathit{f}_{i,\ell}\Longleftrightarrow$
$\left(\bigvee_{i^{\prime}\in\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow i,\ell}\rrbracket}\hskip-15.00002pt\mathit{ref}^{\mathrm{A}}_{i^{\prime}\leftarrow i,\ell}\right)\vee\left(\bigvee_{i^{\prime}\in\llbracket\mathit{ref}^{\mathrm{B}}_{\circ\leftarrow i,\ell}\rrbracket}\hskip-15.00002pt\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow i,\ell}\right)\vee\left(\bigvee_{(i^{\prime},\ell^{\prime})\in\llbracket\mathit{ref}^{\mathrm{C}}_{\circ,\circ\leftarrow i,\ell}\rrbracket}\hskip-23.00006pt\mathit{ref}^{\mathrm{C}}_{i^{\prime},\ell^{\prime}\leftarrow i,\ell}\right).$ (3)
@@
$\forall(i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\cdot\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{B}}_{\cdot\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{C}}_{\cdot,\cdot\leftarrow\circ,\circ}\rrbracket:$
@@
Constraint 4 for a fixed $(i,\ell)$ can be encoded in size of $\left|\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow i,\ell}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{B}}_{\circ\leftarrow i,\ell}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{C}}_{\circ,\circ\leftarrow i,\ell}\rrbracket\right|=O(n^{2})$ using an efficient encoding for this kind of “at-most” constraints [13].
@@
$\forall(i^{\prime},\ell^{\prime},i)\in\llbracket\mathit{dref}_{\circ,\circ\leftarrow\circ}\rrbracket:\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}\Longleftrightarrow$
$\mathit{ref}^{\mathrm{A}}_{i^{\prime}\leftarrow i,\ell^{\prime}}\vee\left(\bigvee_{\ell\in\llbracket\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow i,\circ}\rrbracket}\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow i,\ell}\right)\vee\left(\bigvee_{\ell\in\llbracket\mathit{ref}^{\mathrm{C}}_{i^{\prime},\ell^{\prime}\leftarrow i,\circ}\rrbracket}\mathit{ref}^{\mathrm{C}}_{i^{\prime},\ell^{\prime}\leftarrow i,\ell}\right).$ (5)
We let $\mathit{q}_{i^{\prime},\ell}$ summarize the information on whether $T[i^{\prime}..i^{\prime}+\ell^{\prime})$ corresponds to a node implied by $\mathit{ref}^{\mathrm{A}}_{\cdot\leftarrow\cdot,\cdot}$, $\mathit{ref}^{\mathrm{B}}_{\cdot\leftarrow\cdot,\cdot}$ and $\mathit{ref}^{\mathrm{C}}_{\cdot,\cdot\leftarrow\cdot,\cdot}$:
$\forall(i^{\prime},\ell^{\prime})\in\llbracket\mathit{q}_{\circ,\circ}\rrbracket:\mathit{q}_{i^{\prime},\ell^{\prime}}\Longleftrightarrow\left(\bigvee_{i\in\llbracket\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}\rrbracket}\hskip-18.99995pt\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}\right)\vee\left(\bigvee_{\begin{subarray}{c}(i,\ell)\in\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow\circ,\circ}\\ \mbox{with }\ell^{\prime}=\ell+i-i^{\prime}\end{subarray}}\hskip-18.99995pt\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow i,\ell}\right).$ (6)
@@
The set of intervals indicated by $\mathit{q}_{\cdot,\cdot}$ must be compatible with tree structures. In other words, for any two substrings $T[i^{\prime}_{1}..i^{\prime}_{1}+\ell^{\prime}_{1})$ and $T[i^{\prime}_{2}..i^{\prime}_{2}+\ell^{\prime}_{2})$ with $i^{\prime}_{1}<i^{\prime}_{2}<i^{\prime}_{1}+\ell^{\prime}_{1}<i^{\prime}_{2}+\ell^{\prime}_{2}$, at most one of $T[i^{\prime}_{1}..i^{\prime}_{1}+\ell^{\prime}_{1})$ and $T[i^{\prime}_{2}..i^{\prime}_{2}+\ell^{\prime}_{2})$ can correspond to a node of the grammar tree. Thus, we require that
@@
$\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}\Longrightarrow\forall d\in[1..n](\lnot\mathit{depth}_{i,1,d}\Longrightarrow\lnot\mathit{depth}_{i^{\prime},\ell^{\prime},d-1}).$ (13)
@@
Correctness of SAT formulation: We prove the correctness of our SAT formulation presented in this section for ICS-factorizations. On the one hand, given an internal collage system, it is clear that a truth assignment to Boolean variables based on the definition will satisfy all constraints. On the other hand, given $T$ and a truth assignment satisfying the constraints, we can deduce a factorization to satisfy the conditions (i), (ii) and (iii) of Lemma˜5. First, the truth assignments of $\mathit{p}_{i}$ and Constraint (1) give a factorization of $T$ where we regard $T[i..i+\ell)$ as a factor if and only if $\mathit{f}_{i,\ell}=1$. If $\mathit{f}_{i,\ell}=1$ with $\ell>1$, the type of factor $T[i..i+\ell)$ and its reference to satisfy condition (i) are specified by the truth assignments of $\mathit{ref}^{\mathrm{A}}_{\cdot\leftarrow\cdot,\cdot}$, $\mathit{ref}^{\mathrm{B}}_{\cdot\leftarrow\cdot,\cdot}$ and $\mathit{ref}^{\mathrm{B}}_{\cdot\leftarrow\cdot,\cdot}{\cdot}$, which are suitably controlled by Constraints (2), (3) and (4). By Constraint (5), the information about the starting position of a factor and its reference is collected in variables $\mathit{dref}_{\cdot,\cdot\leftarrow\cdot}$. By Constraint (6), the truth assignments of $\mathit{q}_{\cdot,\cdot}$ give the set $I$ of intervals of Lemma˜5. The starting position of every interval of $I$ must be the starting position of an ICS-factor, which is enforced by Condition (7). Constraint (8) ensures that $I$ is compatible with tree structures to satisfy condition (ii). Finally, condition (iii) is ensured by Constraints (9), (10), (11), (12) and (13) using $\mathit{dref}_{\cdot,\cdot\leftarrow\cdot}$ and $\mathit{depth}_{\cdot,\cdot,\cdot}$.
@@
MAX-SAT formulation: By Proposition˜2, the size $m$ of an internal collage system corresponding to an ICT-factorization is $m=h+m_{\mathsf{tr}}+\sigma-1$, where $h$ is the number of factors, $m_{\mathsf{tr}}$ is the number of truncation rules, and $\sigma$ is the number of distinct characters in $T$. Since $\sigma$ is fixed, $m$ is minimized if $h+m_{\mathsf{tr}}$ is minimized. This can be encoded in MAX-SAT by introducing the set of soft clauses consisting of $\lnot\mathit{p}_{\cdot}$ and $\lnot\mathit{ref}^{\mathrm{C}}_{\cdot,\cdot\leftarrow\cdot,\cdot}$ because satisfying these soft clauses as many as possible leads to minimizing $h+m_{\mathsf{tr}}$.
@@
Size analysis: In total, we have $O(n^{4})$ Boolean variables dominated by $\mathit{ref}^{\mathrm{C}}_{\cdot,\cdot\leftarrow\cdot,\cdot}$. The total size of the resulting MAX-SAT formulation is $O(n^{4})$, dominated by Constraints (3), (4), (5), (6), (8), (12) and (13) where there are four free parameters each of which independently takes $O(n)$ different values. ∎
Acknowledgements
Tomohiro I was supported by JSPS KAKENHI (Grant Numbers JP24K02899) and JST AIP Acceleration Research JPMJCR24U4, Japan.

### v2

@@ 4 MAX-SAT formulation to compute $\hat{c}(T)$
  • $\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}$ for $(i^{\prime},i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow\circ,\circ}\rrbracket$ with $\ell^{\prime}=\ell$, $(i^{\prime},i)\in\llbracket\mathit{ref}^{\mathrm{B}}_{\circ\leftarrow\circ,\bullet}\rrbracket$ with $\ell^{\prime}=i-i^{\prime}$, or $(i^{\prime},\ell^{\prime},i)\in\llbracket\mathit{ref}^{\mathrm{C}}_{\circ,\circ\leftarrow\circ,\bullet}\rrbracket$: $\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}=1$ iff there is an ICS-factor that starts at $i$ and refers to $T[i^{\prime}..i^{\prime}+\ell^{\prime})$.
  • $\mathit{q}_{i^{\prime},\ell^{\prime}}$ for $(i^{\prime},\ell^{\prime})\in\llbracket\mathit{dref}_{\circ,\circ\leftarrow\bullet}\rrbracket$ or $(i^{\prime},i,\ell)\in\llbracket\mathit{ref}^{\mathrm{B}}_{\circ\leftarrow\circ,\circ}\rrbracket$ with $\ell^{\prime}=\ell+i-i^{\prime}$: $\mathit{q}_{i^{\prime},\ell^{\prime}}=1$ iff $i^{\prime}..i^{\prime}+\ell^{\prime})$ is in the set $I$ defined in Lemma˜4.2.
@@
$\forall(i,\ell)\in\llbracket\mathit{f}_{\circ,\circ}\rrbracket\setminus(\llbracket\mathit{ref}^{\mathrm{A}}_{\bullet\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{B}}_{\bullet\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{C}}_{\bullet,\bullet\leftarrow\circ,\circ}\rrbracket)\mbox{ with }l>1:\lnot\mathit{f}_{i,l}.$ (2)
@@
$\forall(i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\bullet\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{B}}_{\bullet\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{C}}_{\bullet,\bullet\leftarrow\circ,\circ}\rrbracket:$
$\mathit{f}_{i,\ell}\Longleftrightarrow\left(\bigvee_{i^{\prime}\in\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow i,\ell}\rrbracket}\hskip-15.00002pt\mathit{ref}^{\mathrm{A}}_{i^{\prime}\leftarrow i,\ell}\right)\vee\left(\bigvee_{i^{\prime}\in\llbracket\mathit{ref}^{\mathrm{B}}_{\circ\leftarrow i,\ell}\rrbracket}\hskip-15.00002pt\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow i,\ell}\right)\vee\left(\bigvee_{(i^{\prime},\ell^{\prime})\in\llbracket\mathit{ref}^{\mathrm{C}}_{\circ,\circ\leftarrow i,\ell}\rrbracket}\hskip-23.00006pt\mathit{ref}^{\mathrm{C}}_{i^{\prime},\ell^{\prime}\leftarrow i,\ell}\right).$ (3)
@@
$\forall(i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\bullet\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{B}}_{\bullet\leftarrow\circ,\circ}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{C}}_{\bullet,\bullet\leftarrow\circ,\circ}\rrbracket:$
@@
Constraint 4 for a fixed $(i,\ell)$ can be encoded in size of $\left|\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow i,\ell}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{B}}_{\circ\leftarrow i,\ell}\rrbracket\cup\llbracket\mathit{ref}^{\mathrm{C}}_{\circ,\circ\leftarrow i,\ell}\rrbracket\right|=O(n^{2})$ using an efficient encoding for this kind of “at-most” constraints [2005Sinz_TowarOptimCnfEncodOf_CP].
@@
$\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}\Longleftrightarrow\mathit{ref}^{\mathrm{A}}_{i^{\prime}\leftarrow i,\ell^{\prime}}\vee\left(\bigvee_{\ell\in\llbracket\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow i,\circ}\rrbracket}\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow i,\ell}\right)\vee\left(\bigvee_{\ell\in\llbracket\mathit{ref}^{\mathrm{C}}_{i^{\prime},\ell^{\prime}\leftarrow i,\circ}\rrbracket}\mathit{ref}^{\mathrm{C}}_{i^{\prime},\ell^{\prime}\leftarrow i,\ell}\right).$ (5)
We let $\mathit{q}_{i^{\prime},\ell}$ summarize the information on whether $T[i^{\prime}..i^{\prime}+\ell^{\prime})$ corresponds to a node implied by $\mathit{ref}^{\mathrm{A}}_{\bullet\leftarrow\bullet,\bullet}$, $\mathit{ref}^{\mathrm{B}}_{\bullet\leftarrow\bullet,\bullet}$ and $\mathit{ref}^{\mathrm{C}}_{\bullet,\bullet\leftarrow\bullet,\bullet}$:
$\forall(i^{\prime},\ell^{\prime})\in\llbracket\mathit{q}_{\circ,\circ}\rrbracket:\mathit{q}_{i^{\prime},\ell^{\prime}}\Longleftrightarrow\left(\bigvee_{i\in\llbracket\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow\circ}\rrbracket}\hskip-18.99995pt\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}\right)\vee\left(\bigvee_{\begin{subarray}{c}(i,\ell)\in\llbracket\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow\circ,\circ}\rrbracket\\ \mbox{with }\ell^{\prime}=\ell+i-i^{\prime}\end{subarray}}\hskip-18.99995pt\mathit{ref}^{\mathrm{B}}_{i^{\prime}\leftarrow i,\ell}\right).$ (6)
@@
If $\mathit{q}_{i^{\prime},\ell^{\prime}}$ is true, then $T[i^{\prime}..i^{\prime}+\ell^{\prime})$ must not be a type-B factor:
$\forall(i^{\prime},\ell^{\prime})\in\llbracket\mathit{q}_{\circ,\circ}\rrbracket,\forall i^{\prime\prime}\in\llbracket\mathit{ref}^{\mathrm{B}}_{\circ\leftarrow i^{\prime},\ell^{\prime}}\rrbracket:\mathit{q}_{i^{\prime},\ell^{\prime}}\Longrightarrow\lnot\mathit{ref}^{\mathrm{B}}_{i^{\prime\prime}\leftarrow i^{\prime},\ell^{\prime}}.$ (8)
The set of intervals indicated by $\mathit{q}_{\bullet,\bullet}$ must be compatible with tree structures. In other words, for any two substrings $T[i^{\prime}_{1}..i^{\prime}_{1}+\ell^{\prime}_{1})$ and $T[i^{\prime}_{2}..i^{\prime}_{2}+\ell^{\prime}_{2})$ with $i^{\prime}_{1}<i^{\prime}_{2}<i^{\prime}_{1}+\ell^{\prime}_{1}<i^{\prime}_{2}+\ell^{\prime}_{2}$, at most one of $T[i^{\prime}_{1}..i^{\prime}_{1}+\ell^{\prime}_{1})$ and $T[i^{\prime}_{2}..i^{\prime}_{2}+\ell^{\prime}_{2})$ can correspond to a node of the grammar tree. Thus, we require that
@@
$\forall(i^{\prime},\ell^{\prime},i)\in\llbracket\mathit{dref}_{\circ,\circ\leftarrow\circ}\rrbracket:\mathit{dref}_{i^{\prime},\ell^{\prime}\leftarrow i}\Longrightarrow\forall d\in[1..n](\lnot\mathit{depth}_{i,1,d}\Longrightarrow\lnot\mathit{depth}_{i^{\prime},\ell^{\prime},d-1}).$ (14)
@@
Correctness of SAT formulation: We prove the correctness of our SAT formulation presented in this section for ICS-factorizations. On the one hand, given an internal collage system, it is clear that a truth assignment to Boolean variables based on the definition will satisfy all constraints. On the other hand, given $T$ and a truth assignment satisfying the constraints, we can deduce a factorization to satisfy the conditions (i), (ii), (iii) and (iv) of Lemma˜4.2. First, the truth assignments of $\mathit{p}_{i}$ and Constraint (1) give a factorization of $T$ where we regard $T[i..i+\ell)$ as a factor if and only if $\mathit{f}_{i,\ell}=1$. If $\mathit{f}_{i,\ell}=1$ with $\ell>1$, the type of factor $T[i..i+\ell)$ and its reference to satisfy condition (i) are specified by the truth assignments of $\mathit{ref}^{\mathrm{A}}_{\bullet\leftarrow\bullet,\bullet}$, $\mathit{ref}^{\mathrm{B}}_{\bullet\leftarrow\bullet,\bullet}$ and $\mathit{ref}^{\mathrm{C}}_{\bullet,\bullet\leftarrow\bullet,\bullet}$, which are suitably controlled by Constraints (2), (3) and (4). By Constraint (5), the information about the starting position of a factor and its reference is collected in variables $\mathit{dref}_{\bullet,\bullet\leftarrow\bullet}$. By Constraint (6), the truth assignments of $\mathit{q}_{\bullet,\bullet}$ give the set $I$ of intervals of Lemma˜4.2. The starting position of every interval of $I$ must be the starting position of an ICS-factor, which is enforced by Constraint (7). Condition (iii) is fulfilled by Constraint (8). Constraint (9) ensures that $I$ is compatible with tree structures to satisfy condition (ii). Finally, condition (iv) is ensured by Constraints (10), (11), (12), (13) and (14) using $\mathit{dref}_{\bullet,\bullet\leftarrow\bullet}$ and $\mathit{depth}_{\bullet,\bullet,\bullet}$.
@@
MAX-SAT formulation: By Proposition˜2.1, the size $m$ of an internal collage system corresponding to an ICS-factorization is $m=h+m_{\mathsf{tr}}+\sigma-1$, where $h$ is the number of factors, $m_{\mathsf{tr}}$ is the number of truncation rules, and $\sigma$ is the number of distinct characters in $T$. Since $\sigma$ is fixed, $m$ is minimized if $h+m_{\mathsf{tr}}$ is minimized. This can be encoded in MAX-SAT by introducing the set of soft clauses consisting of $\lnot\mathit{p}_{\bullet}$ and $\lnot\mathit{ref}^{\mathrm{C}}_{\bullet,\bullet\leftarrow\bullet,\bullet}$ because satisfying these soft clauses as many as possible leads to minimizing $h+m_{\mathsf{tr}}$.
@@
Size analysis: In total, we have $O(n^{4})$ Boolean variables dominated by $\mathit{ref}^{\mathrm{C}}_{\bullet,\bullet\leftarrow\bullet,\bullet}$. The total size of the resulting MAX-SAT formulation is $O(n^{4})$, dominated by Constraints (3), (4), (5), (6), (9), (13) and (14) where there are four free parameters each of which take $O(n)$ different values.
5 Conclusions and future work
In this paper, we proposed an $O(m^{2})$-time algorithm to convert a collage system of size $m$ for a string $T$ to an internal collage system of size $\leq 9m$. As a consequence, we obtained new bounds $\hat{c}(T)=\Theta(c(T))$ and $b(T)=O(c(T))$ on the smallest size of collage systems. We also give a MAX-SAT formulation to compute $\hat{c}(T)$ for a given $T$.
Future work includes the following directions:
  • Improve the current (probably loose) upper bound $9m$ for the size of converted internal collage system or give a nontrivial lower bound.
  • Improve $O(m^{2})$ running time for the conversion algorithm.
  • Implement our MAX-SAT formulation and evaluate how well it scales compared to other formulations for SLPs [2022BannaiGIKKN_ComputNpHardRepetMeasur_ESA] and RLSLPs [2024KawamotoIKB_HardnOfSmallRlslpAnd_DCC].
  • Consider a formulation to compute $c(T)$. Here, a new technique would be required because the current approach, inherited from [2022BannaiGIKKN_ComputNpHardRepetMeasur_ESA, 2024KawamotoIKB_HardnOfSmallRlslpAnd_DCC], depends on the fact that $\hat{c}(T)$ is measured through the ICS-factorization of $T$, but it does not hold for general collage systems that have “external” references.
  • Prove or disprove that $z(T)/c(T)\in O(1)$, which is left unknown in the relation between compressibility measures presented in [2023KociumakaNP_TowarDefinComprMeasurFor, Fig. 1].